it takes a painter A 3 hrs to paint a certain area of a hhouse.it takes painter B 5 hrs to do the same job.how long would it take them,working together,to do the painting job?
Answer by checkley71(8403)

3*5/(3+5)=15/8=1.875 HOURS WHEN THEY WORK TOGETHER.